Determine the energy change associated with the transition from n=2 to n=5 in the hydrogen atom. A) -2.18 × 10-19 J B) +6.54 × 10-19 J C) +4.58 × 10-19 J D) -1.53 × 10-19 J E) +3.76 × 10-19 J
Answer #1

n1 = 2 and n2 = 5

Planck's constant, h = 6.626 x 10-34 Jsec

Velocity of light, C = 3 x 108 m/sec

Rydberg constant, R = 1.0974 x 107 /sec

Energy change associated with the transition, \DeltaE = -Rhc((1/n2)2 - (1/n1)2)

Rhc = 6.626 x 10-34 Jsec x 3 x 108 m/sec x 1.0974 x 107 /m sec = 21.8 x 10-19 J

\DeltaE = -Rhc((1/n2)2 - (1/n1)2) = - 21.8 x 10-19 J (1/25))-(1/4))

= - 21.8 x 10-19 J (4-25/100)

= -21.8 x 10-19 J (-0.21)

= + 4.58 x 10-19 J (OPTION C IS CORRECT)